# The heat flow on glued spaces with varying dimension

### v1

@@ 4 Generalizations
With $\omega_{i}$ satisfying the Muckenhoupt type condition, this yields
$\text{Cap}_{2,\mu_{i}}(L,L_{R})$ $\lesssim\left(R^{2}\mu_{i}(L_{R})^{-1}+2(n_{i}-k-1)\int\limits_{0}^{R}\rho\mu_{i}^{-1}(L_{\rho})\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Next, we prove the iteration procedure for more than two manifolds. We are in the setting of $(X,d,\mu)$ where $X$ is given as the connected union of embedded Euclidean manifolds of varying dimension glued at smooth submanifolds:
$X:=\bigcup\limits_{i=1}^{m}M_{i}.$
Here, the manifolds $M_{i}$ are compact, connected, oriented with smooth boundary and of dimension $n_{i}$. We require that their pairwise intersections $L_{ij}$ are disjoint to each other, lie in the inner of $M_{i},M_{j}$ and are transversal. Additionally, they have to satisfy $\mu(L_{ij})=0$ and that $X$ is path-wise connected. The transversality does not need to be satisfied in the whole space that $X$ is contained in, a local transversality is suitable, i.e. there exists a neighborhood manifold $Y$ containing $\left(L_{ij}\right)_{\varepsilon}$ for some $\varepsilon>0$ s.t. $M_{i}$ and $M_{j}$ intersect transversal in $Y$. We denote the dimension of $L_{ij}$ by $k_{ij}$.
The measure
$\mu:=\sum\limits_{i=1}^{m}\omega_{i}\cdot\text{vol}_{M_{i}}$
and Dirichlet form
$\operatorname*{\mathcal{E}}(f):=\int\limits_{X}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu:=\sum\limits_{i=1}^{m}\int\limits_{M_{i}}|\nabla^{M_{i}}f|^{2}\mathop{}\!\mathrm{d}\mu_{i}$
are defined analogously to before. In particular, $\omega_{i}\in A_{2}(M_{i}),\omega_{i}>0$. Again, we can choose $d$ to be the glued distance. Here, $\operatorname*{\mathcal{E}}$ is strictly local and regular and $\mu$ has to satisfy a $N-$doubling condition with $N\in L^{1}_{loc}$.
Additionally, assume that there are neighborhoods of $L_{ij}$ s.t. $\mu_{j}$ is bounded from above and below and that $k_{ij}+1=n_{j}<n_{i}$.
@@
In this setting, we can use the same proofs as before as the calculations for the capacities are done locally around the intersections and the other proofs are done analogously for a finite number of manifolds. I.e. we still can use that $(X,d)$ is a metric doubling space to prove the compact embedding and with the connectedness and capacity conditions, the ergodicity and irreducibility follow analogously.
@@ 5 Outlook
In this section, we will give the needed notations and definitions to define a non-local perimeter functional, the heat excess $E_{h}$. An interesting question is whether this functional converges to the perimeter as $h$ goes to zero. This proven in the context of Euclidean space (e.g. [19]) and smooth manifolds ([49]) but unknown in our setting or general metric measure spaces. In this context, it is known that the limit leads to an equivalent characterization of BV under assumptions on the measure, c.f. [1, 2, 3, 52, 58, 59].
We define the perimeter functional for a function $f\in L^{1}(X,\mu)$ as (in accordance with [52]):
$\left\lVert Df\right\rVert(U):=\liminf\left\{\left\lVert\nabla f_{n}\right\rVert(U)\left|f_{n}\to f\text{ in }L^{1},f_{n}\in\operatorname*{Lip}\right.\right\}.$
In this note, we use $\nabla f$ if the density exists as a function (absolutely continuous) and $Df$ for a general derivative or in this case the perimeter / total variation.
@@
$E_{h}(f):=\int\limits_{X}S_{h}(x,y)|f(x)-f(\cdot)|(x)\mathop{}\!\mathrm{d}\mu(x).$
@@
$E_{h}(f):=\int\limits_{X}\int\limits_{X}p_{h}(x,y)|f(x)-f(y)|\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
@@
$\mathcal{E}_{h}(f)=\int\limits_{X}\chi_{U^{C}}S_{h}\chi_{U}\mathop{}\!\mathrm{d}\mu(x).$
This can also be seen as a Korevaar-Schoen type energy. Thus, the convergence is related to the Bourgain-Brezis-Mironescu formula, c.f. [31].
An interesting question is whether the following conjecture holds true in general or even whether it is true in the setting of glued manifolds with $A_{1}$ weights and doubling measure.
@@
Let $(X,d,\mu)$ be a complete metric space with $(\operatorname*{\mathcal{E}},D(\operatorname*{\mathcal{E}}))$ a strictly local, regular Dirichlet form. Moreover, assume that $\mu$ satisfies a $1-$Poincare inequality and is doubling. Then, for $f\in BV(X)$:
$E_{h}(f)\overset{L^{1}-\Gamma}{\longrightarrow}\left\lVert Df\right\rVert(X).$
Given this convergence, we can construct a classifier by considering a minimizing movement scheme similar to the thresholding scheme (MBO-scheme) to receive a fast iterative algorithm to classify high-dimensional data.
In proving the $\Gamma-$convergence of the heat excess $E_{h}(f)\overset{\Gamma}{\to}\left\lVert Df\right\rVert$ in $L^{1}$, it is necessary for $\mu$ to satisfy a $1-$Poincare inequality (or similar condition on the measure $\mu$).
Definition 5.1 ($p-$Poincare inequality).
We say that $(X,d,\mu)$ satisfies a $(q,p)-$Poincare inequality if, for each $f\in\operatorname*{Lip}$, there exists $C>0$ and a dilation factor $\lambda\geq 1$ s.t. for each ball $B$:
@@
If $p=q$, we will call this condition also a $p-$Poincare inequality.
@@
If $X$ satisfies a $(q,p)-$Poincare inequality then the same is true for $1\leq q^{\prime}\leq q$ but we also can recover a $(p,p)-$Poincare inequality in doubling spaces. Moreover, if $X$ satisfies a $(1,p)-$Poincare inequality, it satisfies a $(1,p^{\prime})-$Poincare inequality for each $1\leq p^{\prime}\leq p$.
In the setting of glued manifolds, for a $1-$Poincare inequality to hold on the whole space, it is required to hold on the components and a $1-$capacity condition needs to be satisfied.
@@
The relative $p-\mu-$capacity is defined on compact sets and extended to measurable sets in the usual way:
$\text{Cap}_{p,\mu}(K,\Omega):=\inf\left\{\left.\int\limits_{\Omega}|\nabla u|^{p}\mathop{}\!\mathrm{d}\mu\right|u\in\operatorname*{Lip}\cap C_{c}(\Omega),u\geq\mathds{1}_{K}\right\}.$
Moreover, with $1-$Muckenhoupt weights, the $1-$Poincare inequality is satisfied on the individual manifolds.
@@
$\fint\limits_{B}\omega_{i}\mathop{}\!\mathrm{d}\text{vol}_{M_{i}}\cdot\left(\fint\limits_{B}\frac{1}{\omega_{i}^{\frac{1}{p-1}}}\mathop{}\!\mathrm{d}\text{vol}_{M_{i}}\right)^{p-1}\lesssim 1$
@@
$\fint\limits_{B}\omega_{i}\mathop{}\!\mathrm{d}\text{vol}_{M_{i}}\simeq\operatorname*{ess\,inf}_{B}\omega.$
For the future, an aim is to investigate the $\Gamma-$ or Mosco–convergence of $E_{h}$ to $\left\lVert Df\right\rVert$. To this end, different inequalities as well as bounds for the heat kernel are interesting. It is notable that one has to consider $A_{1}$ weights and $1-$Poincare inequalities instead of the $L^{2}$ variants for this. Moreover, to carry over these inequalities from the components to the whole space, one needs more refined geometric assumptions on the intersection related to $\text{Cap}_{1,\mu}$.
Another aim is to investigate more refined properties of the heat flow and associated Brownian motion. Additionally, for the iteration and general gluing, even at already existing intersections, it is interesting to investigate the assumptions on the glued space, i.e. whether the inequalities like the 2-Poincare inequality are still satisfied globally.
@@ 6 Related work
In this work, we consider the setting of glued manifolds with weights. This setting originates from manifold learning and examples of similar approaches were studied before in the works we list next. In [15, 54] the space $\mathbb{R}^{2}\times\mathbb{R}_{+}$ with a collapsed center is analyzed and an expansion of the heat kernel and Brownian motion is computed. Similarly, [50, 51] study so-called distorted Brownian motion in $\mathbb{R}^{3}\times\mathbb{R}_{+}$ where the connection is done via a density that fulfills our conditions. Using the same setting, in [54], Ooi extends the results of [15] to the case of two arbitrary $\mathbb{R}^{d_{1}},\mathbb{R}^{d_{2}}$ meeting at one collapsed point with explicit heat kernel bounds. The works [11, 16] study bow ties as special structures and how Poincare inequalities on them can be obtained from capacity conditions. In [55], glued spaces with constant dimension are analyzed. A major difference is that only manifolds of the same dimension are considered and they don’t posses a density. Finally, in [28], the authors looks at heat kernels on manifolds with ends.
The theory of Muckenhoupt weights and related study of them can be found in a various papers and books. The main sources for this work were [40], [36, Chapter 2 and 15] and [61, Chapter 5] where they are introduced and studied together with the related capacities. For the setting of metric measure spaces, one can see [46] or in the context of non-doubling measures [42].
In their series of four paper [6, 7, 5, 8] the authors investigate, among other things, Muckenhoupt weights on manifolds, where the appendix in [7] is especially noteworthy.
@@
The theory of heat flows and kernels with their estimates is extensive and contained in a lot of fields. In [26], Grigor’yan and co-authors provide a condition for the reverse doubling condition on manifolds and prove connections between measure doubling conditions and heat kernel estimates. In [21], he provides a general theory of upper and lower bounds of heat kernels and how to achieve these in a general setting. Other estimates on heat kernels by Grigor’yan are proven in [22, 23, 24, 25, 27].
@@
In [63] and [62], Sturm shows heat kernel bounds under a doubling assumption on the measure. For the bounds on heat kernels, [14] provides theorems which imply an upper bound given a Sobolev inequality of Nash type. This Nash inequality on the manifolds can be proven using [45] where a proof is given for weighted manifolds which fit our setting. Other Nash inequalities can be found in [9, 39].
@@
In [44], a heat flow and Brownian motion are constructed from Dirichlet forms which arise as limits of discrete forms. For the expansion of the heat kernel on manifolds with smooth weights, a parametrix construction is needed which is done in [57]. Even in a more general setting, the heat kernel localizes which can be seen in locality theorems and a comparison result in [56]. In [10], the Lipschitz continuity of heat kernels in metric measure spaces is investigated. More bounds on the heat kernel can be found in [12, 17].
@@
In [33], Hajłasz defines different notions of Sobolev spaces and in [34], he proves for instance Sobolev embeddings of them. A short summary of Hajłasz Sobolev-spaces is given in [32] and [32]. Generalizations of embeddings of Sobolev spaces can be found in [30, 37]. Here, the former analyzes the setting of embeddings in weighted Sobolev spaces whereas the latter gives a generalization of the Rellich-Kondrachov embedding for the Hajłasz Sobolev-space. This proof uses a similar criteria to our $N$-doubling condition.
@@
Functions of bounded variations possess numerous different characterizations in metric spaces. The works [53] and [13] give an overview and characterizations of different definitions. Moreover, in [4] a BMO characterization of sets of finite perimeter is given.
@@
The convergence of the heat excess was already studied in different cases. The most relevant ones to this setting are [1, 2, 3, 19, 49, 52, 58, 59]. First, in [19] the convergence in Euclidean space for the multi-phase mean curvature flow setting is shown. In [49], the $\Gamma$-convergence was proven in the setting of smooth manifolds. [1] as well as [2], [3] and [59] use a weak Bakry Emery condition in the setting of metric measure spaces with a $2$-Poincare inequality to establish equivalence of the limit with the BV seminorm. In comparison, [52] and [58] prove similar results using a $1$-Poincare inequality. The latter is in the setting of smooth manifolds where pointwise convergence is shown.
@@
Similar to this question, is the one after a Bourgain-Brezis-Mironescu formula in metric measure spaces, i.e. whether a non-local kernel perimeter converges to the perimeter. This was investigated in [18, 31, 35, 43, 47, 48].

### v2

@@ 4 Generalizations
With $\omega_{2}$ satisfying the Muckenhoupt type condition, this yields
${\text{Cap}}_{2,\mu_{2}}(L,L_{R})$ $\lesssim\left(R^{2}\mu_{2}(L_{R})^{-1}+2(n_{2}-k-1)\int\limits_{0}^{R}\rho\mu_{2}(L_{\rho})^{-1}\mathop{}\!\mathrm{d}\rho\right)^{-1}.$
Now, we can conclude that the capacities are positive under the assumptions of Theorem 4.1. By assumption and since $R>0$ is fixed, we have
$N(R)R^{2}\mu_{2}(L_{R})^{-1},R^{2}\mu_{2}(L_{R})^{-1}<\infty.$
Additionally, we assumed $\int\limits_{0}^{R}\frac{N(\rho)\rho}{\mu_{2}(L_{\rho})}\mathop{}\!\mathrm{d}\rho,\int\limits_{0}^{R}\frac{\rho}{\mu_{2}(L_{\rho})}\mathop{}\!\mathrm{d}\rho<\infty$ and thus by our lower bound, we get
${\text{Cap}}_{2,\mu_{1}}(L,L_{R}),{\text{Cap}}_{2,\mu_{2}}(L,L_{R})>0.$
Example 4.2.
The following example demonstrates the effective dimension condition when the weight is locally a rational function around the intersection. Additionally, it shows that the $L$-Muckenhoupt type condition is reasonable.
We consider a submanifold $L$ in a manifold $M$ with $n\coloneqq\dim M$ and $k\coloneqq\dim L\neq n-1$ with weight $\omega$ which locally around $L$ has the form
$\omega(x)\coloneqq\frac{1}{\operatorname*{dist}(x,L)^{\alpha}}.$
Then locally,
$\frac{\mathcal{L}^{n}}{\omega}(L_{r})=\int\limits_{L_{r}}\frac{1}{\omega}\mathop{}\!\mathrm{d}x$ $\simeq\begin{cases}\frac{r^{n-k+\alpha}}{n-k+\alpha}&\text{if }\alpha>-(n-k),\\ \infty&\text{else.}\end{cases}$
$\int\limits_{0}^{R}\rho\mu_{i}(L_{\rho})^{-1}\mathop{}\!\mathrm{d}\rho$ $=\int\limits_{0}^{R}\rho^{1-n+k+\alpha}\mathop{}\!\mathrm{d}\rho<\infty,$
i.e., $\alpha>n-k-2$, which coincides with our intuition that the difference of $k$ and effective dimension minus one, $n-\alpha-1$, has to be less than one.
Next, we describe the iteration procedure for more than two manifolds. We define the Dirichlet form analogously to before
$\operatorname*{\mathcal{E}}(f)\coloneqq\int\limits_{X}|\nabla f|^{2}\mathop{}\!\mathrm{d}\mu\coloneqq\sum\limits_{i=1}^{m}\int\limits_{M_{i}}|\nabla^{M_{i}}f|^{2}\mathop{}\!\mathrm{d}\mu_{i}.$
@@
In this setting, we can use the same proofs as before since the calculations for the capacities are done locally around the intersections and the other proofs are done analogously for a finite number of manifolds. Therefore, we still can use that $(X,d)$ is a metric doubling space to prove the compact embedding. We need to use the connectedness to conclude that the kernel again consists only of the constants. Then, the capacity conditions, the ergodicity and irreducibility follow analogously.
Moreover, we can treat other boundary conditions and manifolds in a similar fashion. If we additionally have the existence of a heat kernel with suitable upper and lower exponential bounds, we can use the localization result in metric measure spaces, see [62, Theorem 3.22], and the semigroup property to extend our results to more general settings. This is due to the fact an everywhere positive heat kernel implies ergodicity and irreducibility. In this setting, only the local geometry matters and thus, different boundary conditions can be treated as well.
@@ 5 Outlook
In the following, we give the notation and definitions needed to define a non-local perimeter functional, the heat excess $E_{h}$. An interesting question is whether this functional converges to the perimeter as $h$ goes to zero. The correct notion of convergence for this is the one of $\Gamma$-convergence, cf. Definition 5.1. This convergence is proven in the context of Euclidean space, see [23], and smooth manifolds, see [53], but it is unknown in our setting or for general metric measure spaces. In this context, it is known that the limit leads to an equivalent characterization of $\operatorname*{BV}$ under assumptions on the measure; cf. [1, 2, 3, 4, 5, 56].
We begin with the definition of $\Gamma$-convergence.
Definition 5.1 ($\Gamma$-convergence).
Let $X$ be a topological space and $(F_{n})_{n\in\mathbb{N}}$ be a sequence of functionals $F_{n}:X\to\mathbb{R}\cup\{\infty\}$. Then, we say that $(F_{n})_{n\in\mathbb{N}}$ $\Gamma$-converges to a functional $F:X\to\mathbb{R}\cup\{\infty\}$, written $F_{n}\overset{\Gamma}{\to}F$ if
  • for all $x_{n}\to x,x_{n},x\in X$ we have the liminf inequality
    $\liminf\limits_{n}F_{n}(x_{n})\geq F(x),$
  • for all $x\in X$ there exists a recovery sequence $x_{n}\to x,x_{n}\in X$ such that
    $\lim\limits_{n}F_{n}(x_{n})=F(x).$
Remark.
The second property is often written as the equivalent formulation of a limsup inequality
$\limsup\limits_{n}F_{n}(x_{n})\leq F(x).$
We define the perimeter functional for a function $f\in L^{1}(X,\mu)$ in accordance with [56] as:
$\left\lVert\nabla f\right\rVert(U)\coloneqq\liminf\left\{\left\lVert\nabla f_{n}\right\rVert(U):f_{n}\to f\text{ in }L^{1}(X),f_{n}\in\operatorname*{Lip}(X)\right\}.$
@@
$E_{h}(f)\coloneqq\int\limits_{X}S_{h}|f(x)-f(\cdot)|(x)\mathop{}\!\mathrm{d}\mu(x).$
@@
$E_{h}(f)\coloneqq\int\limits_{X}\int\limits_{X}p_{h}(x,y)|f(x)-f(y)|\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
@@
$E_{h}(f)=\int\limits_{X}\chi_{U^{c}}S_{h}\chi_{U}\mathop{}\!\mathrm{d}\mu(x).$
The heat excess can also be seen as a Korevaar-Schoen type energy. Thus, the convergence is related to the Bourgain-Brezis-Mironescu formula, cf. [25].
An interesting question is whether the following conjecture holds in general or even if it holds in the setting of glued manifolds with $A_{1}$ weights and doubling measure.
@@
Let $(X,d,\mu)$ be a complete metric space with $(\operatorname*{\mathcal{E}},D(\operatorname*{\mathcal{E}}))$ a strictly local regular Dirichlet form. Moreover, assume that $\mu$ satisfies an $1$-Poincaré inequality and is doubling. Then, for $f\in BV(X)$:
$E_{h}(f)\overset{L^{1}-\Gamma}{\longrightarrow}\left\lVert\nabla f\right\rVert(X).$
Given this convergence, we can construct a classifier by considering a minimizing movement scheme similar to the thresholding scheme (MBO-scheme) to receive a fast iterative algorithm to classify high-dimensional data. Schemes like these are analyzed in [70].
In proving the $\Gamma$-convergence of the heat excess $E_{h}(f)\overset{\Gamma}{\to}\left\lVert\nabla f\right\rVert$ in $L^{1}$, it is necessary for $\mu$ to satisfy a $1$-Poincaré inequality (or similar condition on the measure $\mu$).
Definition 5.2 ($p$-Poincaré inequality).
We say that $(X,d,\mu)$ satisfies a $(q,p)$-Poincaré inequality if there exists a constant $c>0$ and a dilation factor $\lambda\geq 1$ such that for all integrable $f$ and for each ball $B$ of radius $r$:
@@
If $p=q$, we call this condition also a $p$-Poincaré inequality.
@@
If $X$ satisfies a $(q,p)$-Poincaré inequality, then we also have a $(q^{\prime},p^{\prime})$-Poincaré inequality for $1\leq q^{\prime}\leq q,p\leq p^{\prime}$. Moreover, we can also recover a $(p,p)$-Poincaré inequality in doubling spaces, cf. [38, Theorem 5.1].
In the setting of glued manifolds, for an inequality $1$-Poincaré to hold throughout the space, it is necessary that the components satisfy the inequality and a $1$-capacity condition must be satisfied at the intersections; cf. [13].
@@
The relative $(p,\mu)$-capacity is defined on compact sets and extended to measurable sets in the usual way:
${\text{Cap}}_{p,\mu}(K,\Omega)\coloneqq\inf\left\{\int\limits_{\Omega}|\nabla u|^{p}\mathop{}\!\mathrm{d}\mu:u\in\operatorname*{Lip}\cap C_{c}(\Omega),u\geq\chi_{K}\right\}.$
Moreover, with $1$-Muckenhoupt weights, the $1$-Poincaré inequality is satisfied in the individual manifolds.
@@
$\fint\limits_{B}\omega\mathop{}\!\mathrm{d}\text{vol}_{M}\cdot\left(\fint\limits_{B}\frac{1}{\omega^{\frac{1}{p-1}}}\mathop{}\!\mathrm{d}\text{vol}_{M}\right)^{p-1}\lesssim 1$
@@
$\fint\limits_{B}\omega\mathop{}\!\mathrm{d}\text{vol}_{M}\simeq\operatorname*{ess\,inf}_{B}\omega.$
Remark.
The spaces of Muckenhoupt weights are nested, i.e., $A_{1}\subseteq A_{p}\subseteq A_{q}$ for $1<p<q$, cf. [8, Proposition A.1].
We can now ask the question of the $\Gamma$- or Mosco-convergence of $E_{h}$ to $\left\lVert\nabla f\right\rVert$. To this end, different inequalities as well as bounds for the heat kernel are interesting. We note that for this one has to consider the $A_{1}$ weights and $1$-Poincaré inequalities instead of the $L^{2}$ variants. Moreover, to carry over these inequalities from the components to the whole space, one needs more refined geometric assumptions on the intersection related to ${\text{Cap}}_{1,\mu}$.
@@ 6 Related work
It is not feasible to give an extensive overview of the heat equation, Brownian motion, heat kernel estimates, and flows on manifolds and metric measure spaces. Thus, we list some of the directly connected more recent works in these areas that are relevant to this paper and our setting.
In this work, we consider the setting of glued manifolds with weights. This setting is motivated by manifold learning, and examples of similar approaches were studied before in the works we list next. There are different approaches that one could follow to connect manifolds. In [19, 59] the space $\mathbb{R}^{2}\times\mathbb{R}_{+}$ with a collapsed center is analyzed and an expansion of the heat kernel and Brownian motion is computed. Similarly, [54, 55] study the so-called distorted Brownian motion in $\mathbb{R}^{3}\times\mathbb{R}_{+}$ where the connection is done via a density that fulfills our conditions. Using the same setting, in [59], Ooi extends the results of [19] to the case of two arbitrary $\mathbb{R}^{n_{1}},\mathbb{R}^{n_{2}}$ meeting at one collapsed point with explicit heat kernel bounds. The works [13, 20] study bow ties as special structures and how Poincaré inequalities on them can be obtained from capacity conditions. In [61], glued spaces with constant dimension are analyzed. A major difference is that only manifolds of the same dimension are considered and they do not possess a density.
In a more recent work than this paper, namely [17], the authors consider Dirichlet energies on intersection manifolds of potentially different dimensions with bounded densities. They derive a formulation that respects the intrinsic dimensions without intrinsic knowledge of them. In accordance with this paper, they obtain the result that the energies separate if the jump in dimension is too high. Furthermore, the energy connects the spaces via a flow condition of the derivative in the case that the manifolds are of the same dimension and intersect in a manifold that has codimension one in them.
The theory of Muckenhoupt weights and related study of them can be found in various papers and books. The main sources for this work were [40, Chapter 2 and 15], [44] and [65, Chapter 5], where they are introduced and studied together with the related capacities. For the setting of metric measure spaces, one can see [50] or, in the context of non-doubling measures, [46].
In their series of four papers [7, 8, 9, 10], the authors investigate, among other things, Muckenhoupt weights on manifolds, where the appendix in [8] is especially noteworthy.
@@
The theory of heat flows and kernels with their estimates is extensive and is contained in many fields. In [30], Grigor’yan and co-authors provide a condition for the reverse doubling condition on manifolds and prove connections between measure doubling conditions and heat kernel estimates. In [26], he provides a general theory of the upper and lower bounds of heat kernels and how to achieve these in a general setting. Other estimates of heat kernels by Grigor’yan are proven in [27, 28, 29, 31, 33].
@@
For the heat kernel, in [67] and [68], Sturm shows bounds under the doubling assumption on the measure. For the bounds on heat kernels, [18] provides theorems that imply an upper bound given a Sobolev inequality of Nash type. This Nash inequality in manifolds can be proven using [49] where a proof is given for weighted manifolds which fit our setting. Other Nash inequalities can be found in [11, 43].
@@
In [48], a heat flow and Brownian motion are constructed from Dirichlet forms which arise as limits of discrete forms. For the expansion of the heat kernel on manifolds with smooth weights, a parametrix construction is needed, which is done in [63]. Even in a more general setting, the heat kernel is localized, which can be seen in the locality theorems and a comparison result in [62]. In [12], the Lipschitz continuity of heat kernels in metric measure spaces is investigated. More bounds on the heat kernel can be found in [14, 21].
Furthermore, in [32], heat kernel bounds on connected manifolds are studied. For the setting, the authors consider same dimensional manifolds that potentially have different scalings in the long-time and large-scale behavior. These are connected through a common set so that a manifold with ends is created for whose heat kernel bounds are derived.
@@
In Section 2, we work with Sobolev spaces that represent the domain of the Dirichlet form and the Laplace-Beltrami operator. In [37], Hajłasz defines different notions of Sobolev spaces, and in [38], he proves, for instance, Sobolev embeddings of them. A short summary of Hajłasz Sobolev-spaces is given in [36]. Generalizations of embeddings of Sobolev spaces can be found in [35, 41]. Here, the former analyzes the setting of embeddings in weighted Sobolev spaces, whereas the latter gives a generalization of the Rellich-Kondrachov embedding for the Hajłasz Sobolev-space. This proof uses a similar criterion to our $N$-doubling condition.
@@
For the classifier and definition perimeter functional, we need to define functions of bounded variations. These possess numerous different characterizations in metric spaces. The works [16] and [57] give an overview and characterization of different definitions. Moreover, in [6] a BMO characterization of sets of finite perimeter is given.
@@
The convergence of the heat excess has already been studied in different cases. The most relevant ones for this setting are [1, 2, 3, 4, 5, 23, 53, 56]. First, in [23] the convergence in Euclidean space for the multiphase mean curvature flow setting is shown. In [53], the $\Gamma$-convergence was proven in the setting of smooth manifolds. [1] as well as [3], [4] and [5] use a weak Bakry-Émery condition in the setting of metric measure spaces with a $2$-Poincaré inequality to establish the equivalence of the limit with the BV seminorm. In comparison, [56] and [2] prove similar results using a $1$-Poincaré inequality. The latter is in the setting of smooth manifolds where pointwise convergence is shown.
@@
Similarly to this question, one can ask for a Bourgain-Brezis-Mironescu formula in metric measure spaces, i.e., whether a nonlocal kernel perimeter converges to the perimeter. This was investigated in [22, 25, 39, 47, 51, 52].